Layoffs looming. American Products plans to lay off 10% of its employees in its aerospace division and 15% of its employees in its agricultural division. If altogether 12% of the 3000 employees in these two division will be laid off, then how many employees are in each division?
Answer by checkley71(8403)

.1X+.15(3000-X)=.12*3000 OR .1X+450-.15X=360 OR -.05=-90 OR X=-90/-.05 OR X=1800 THUS .1*1800+.15(3000-1800)=360 OR 180+.15*1200=360 OR 180+180=360 OR 360=360